cheesesauce:

Which of the following is a linear function? a. y = x3 + 1 b. five over sixx = y − 4 c. y = 8x2 d. y = one over five x

Echofire:

It is a little unclear why some of the numbers are like this. Are there any exponents?

jhonyy9:

like a first step what mean linear function ?

cheesesauce:

@echofire wrote:
It is a little unclear why some of the numbers are like this. Are there any exponents?
Here's the image if it helps

jhonyy9:

a: y = x^3 +1 b. \[\frac{ 5 }{ 6 }x = y - 4\] this i wan to do it

jhonyy9:

@jhonyy9 wrote:
like a first step what mean linear function ?
do you can answer this my question please ?

cheesesauce:

@jhonyy9 wrote:
a: y = x^3 +1 b. \[\frac{ 5 }{ 6 }x = y - 4\] this i wan to do it
My bad, so like this?

jhonyy9:

these posted images are the same - my opinion

cheesesauce:

For anyone out there who has this same question, the answer isn't A. y = x^3 +1 that's all I have to say

Extrinix:

Well a line is \(\sf{y=mx+b}\) So using that, can you figure out which answer is correct?

Extrinix:

I’ll give a hint: \(\sf{+b}\) is optional

cheesesauce:

@extrinix wrote:
I’ll give a hint: \(\sf{+b}\) is optional
y = one over five x ?

Extrinix:

@cheesesauce wrote:
@extrinix wrote:
I’ll give a hint: \(\sf{+b}\) is optional
y = one over five x ?
Does it fit into the \(\sf{y=mx+b}\) equation?

Extrinix:

(That would be yes)